An electromotive force of a cell is .8 5 V and is connected to a load of 4.25Omega , the voltage drops to 2.0 V. Find the current and internal resistance of the cell.



1
Expert's answer
2022-04-19T12:06:01-0400
"I=\frac{\cal {E}}{R+r}"

"V=IR"

Hence, the current

"I=\frac{V}{R}=\frac{2.0}{4.25}=0.47\;\rm A"

and, the internal resistance

"r=\frac{\cal E}{I}-R=8.5/2.0-4.25=2.25\:\Omega"